OpenStudy (anonymous):

is there any technique to plot graphycally any given function without calculation

OpenStudy (asnaseer):

you could use wolframalpha to plot the function. see this as an example: http://www.wolframalpha.com/input/?i=y%3Dx^2%2B3x-8

OpenStudy (amistre64):

the function is a rule that is followed to determine the exact points on the graph. so no, you cannot plot a given function without actually working out the function.

OpenStudy (amistre64):

if you want to sketch a given function that conforms to the elementary functions, then you can eyeball it and use the adjustments to the elementary function to shape it correctly

OpenStudy (anonymous):

Sketch the graph for the function f ( x ) = xe-aX and locate its maximas and minimas ( if any ).

OpenStudy (amistre64):

i cant read the function you wrote ...

OpenStudy (anonymous):

Sketch the graph for the function f ( x ) = xe-aX and locate its maximas and minimas ( if any ).

OpenStudy (amistre64):

f(x)=xe-aX means nothing to me ... it is either badly notated, or it pertains to something more specific in your own material

OpenStudy (anonymous):

f(x)=xexp(-ax2)

OpenStudy (anonymous):

f(x)=x exp(-ax2)

OpenStudy (amistre64):

\[f(x)=x\ exp(-ax^2)\]

OpenStudy (anonymous):

Sketch the graph for the function f(x)=x exp(-ax2) and locate its maximas and minimas ( if any ).

OpenStudy (amistre64):

we can see that this is 0 at x=0 when x= some positive number this follows the same path as e^{-x^2} and when x is negative it follows the same path as -e^{-x^2}

OpenStudy (anonymous):

what is the path of-e^{-x^2} ???????????

OpenStudy (amistre64):